OpenStudy (anonymous):

Could someone help me solve 3x^2 + 5x + 2 = 0 by factoring?

OpenStudy (anonymous):

Any idea about breaking the middle term?

OpenStudy (anonymous):

@KissMyAxe I've never heard of breaking the middle term. I thought the first step would be to multiply 3 times 2 = 6. Rewrite the equation 3x^2 + 3x + 2x + 6.

OpenStudy (anonymous):

The 3x and the 2x were what multiplied to equal 6 and added to equal 5.

OpenStudy (anonymous):

Great! U are doing good :) I was taught that so.... U can factorise it now @jsmusic3

OpenStudy (anonymous):

@KissMyAxe That's where I am getting confused is what the common term is. I tried x(3x+3) and 3(3x+3) That didn't get me back to my equation.

OpenStudy (anonymous):

How about using the quadratic formula?

OpenStudy (anonymous):

@KissMyAxe Where could I find my a, b, and c values in 3x^2 + 3x + 2x + 6?

OpenStudy (anonymous):

YOu don't need to the equation is 3x^2+5x+6=0.

OpenStudy (anonymous):

Okay I'll try, 3x^2+2x+3x+2=0 In the first two terms x is common for both so we take it out 'x(3x+2)' and for the last two 1 is common so '1(3x+2)' The whole thing become like this x(3x+2)+1(3x+2)= (x+1)(3x+2)

OpenStudy (anonymous):

@KissMyAxe I thought that the equation was 3x^2 + 3x + 2x + 6? Does the lat term change back to 2 when you are finding the common term?

OpenStudy (anonymous):

U can write it in both ways, no need to worry! @jsmusic3

OpenStudy (anonymous):

@KissMyAxe So they are interchangeable? It doesn't seem like if I factored (x+1) and (3x+2) I would come out with 3x^2 + 3x + 2x + 6.

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

@KissMyAxe I am sorry, I'm just confused on how they are interchangeable if they don't factor out to the same equation?

OpenStudy (anonymous):

Ahh sorry!!! I didn't read ur last sentence properly! they do come out as 3x^2 + 3x + 2x + 6 @jsmusic3

OpenStudy (anonymous):

@KissMyAxe The last term would come out to 2. since 1 times 2 =2.

OpenStudy (anonymous):

Yes @jsmusic3

OpenStudy (anonymous):

 @KissMyAxe How do you get +6 in the last term if they are interchangeable?